# UniGS: Unified Language-Image-3D Pretraining with Gaussian Splatting

### v1

@@ 1 Introduction
The remarkable success of 2D Image-Text pre-training through modality alignment via contrastive learning (Radford et al., 2021; Sun et al., 2023; Fang et al., 2023; Schuhmann et al., 2022; Qi et al., 2020; Changpinyo et al., 2021; Hong et al., 2021) has recently inspired a line of work pursuing 3D pre-training (Xue et al., 2023a; b; Zeng et al., 2023; Zhang et al., 2023; Liu et al., 2024; Zhang et al., 2022; Huang et al., 2023; Afham et al., 2022). Leveraging diverse large-scale 3D datasets, recent works such as  (Liu et al., 2024; Zhang et al., 2023) expand the traditional 2D task to text-image-3D pertaining by including point clouds as 3D representations, resulting in considerable improvements in 3D zero-shot/open-world object detection, classification and retrieval tasks. While point clouds serve as a natural step towards 3D representations, there are inherent limitations when using them to represent 3D objects. In particular, as illustrated in Fig. 1, point clouds consist of a discrete set of points and thus struggle to accurately capture fine-grained geometric details and surface textures of common objects, limiting the performance of 3D representation learning approaches. Moreover, there exists a noticeable gap between the discrete points and the dense 2D pixels of images, which further hinders the learning of joint multi-modal representations.
@@
However, simply retraining existing multi-modal frameworks like CLIP2 (Zeng et al., 2023) or Uni3D (Zhang et al., 2023) to leverage 3DGS as the 3D representation is not effective. As the spatial connections of 3DGS may be insufficient to capture and express objects due to the fact that 3DGS are not necessarily distributed on the surface of objects.
@@ 2 Related Work
Multi-modal pretraining via contrastive learning. Leveraging multi-modal data to pre-train modality-specific encoders via contrastive alignment has received considerable attention in recent years (Radford et al., 2021; Mu et al., 2022) due to its ability to leverage massive in-the-wild datasets of paired data. While early works have largely been focused on image-text data, promoting research on text-based image manipulation (Patashnik et al., 2021), open vocabulary object detection (Gu et al., 2021; Gao et al., 2022), language grounding (Li et al., 2022) and zero-shot segmentation (Xu et al., 2023), there has been an increasing focus on learning 3D representations lately (Xue et al., 2023a; b; Zeng et al., 2023; Zhang et al., 2023). These approaches largely follow the contrastive learning paradigm by adding a new 3D representation encoder and aligning it with the 2D and/or text modalities. However, due to the difficulty of collecting and constructing 3D representation data, current approaches leverage point clouds. More specially, PointCLIP (Zhang et al., 2022) and CLIP2Point (Huang et al., 2023) extract depth maps from point clouds to obtain image-like data that can be leveraged for contrastive pretraining. Since these depth maps lose plenty of spatial information of the original point cloud data structure, CLIP2 (Zeng et al., 2023) instead learns a 3D encoder directly on point clouds, demonstrating robustness in real-world indoor and outdoor scenarios. However, the flexibility and simplicity of point clouds come at a cost as specific shape and texture information of the object surface is lost, leading to ambiguities. In this work, we, therefore, adopt 3DGS (Kerbl et al., 2023) to replace point clouds as the 3D representation to alleviate this loss of information.
@@ 3.3 Gaussian-Aware Guidance
For better modeling and understanding of the 3D domain represented by 3DGS, we further propose a Gaussian-aware Guidance module. Specifically, this module, as highlighted in the dashed box in Fig. 3a, formulates a dual-branch 3D encoder, where the fundamental encoder $E_{fun}$ leverages the 3D ViT Encoder pre-trained on 3D point clouds from (Zhang et al., 2023) to model the low-level features $f_{fun}$ including spatial and color information of 3DGS, while the advanced encoder $E_{adv}$ additionally models the high-level features $f_{adv}$, the relationship between spatial connections and 3DGS feature. As the number of Gaussian spheres allocated to an object or scene represented by 3DGS varies, a group divider is leveraged to process the input 3DGS into a fixed number of Gaussian spheres. The CNN layers map low-dimensional raw features to a high-dimensional feature space, consistent with Uni3D. Moreover, cross-attention layers are built to extract guidance of embeddings from the fundamental encoder and improve the alignment of the 3DGS features with other domains. We denote the process as $CA$, where
@@ 3.4 Scaling Up 3DGS Backbones
Initializing UniGS. Restricted by the scale of the 3D dataset, directly pre-training each 3D backbone for specific 3D tasks leads to expensive training costs and may suffer from difficulties in convergence or overfitting. To overcome this issue, we follow Uni3D (Zhang et al., 2023) and adopt off-the-shelf pre-trained large models (Sun et al., 2023; Fang et al., 2023; Radford et al., 2021; Caron et al., 2021) with ViT-based structure in the other modalities as the initialization of the 3D backbone to transfer their rich underlying representational abilities to language-image-3D pre-training Different from Uni3D (Zhang et al., 2023), UniGS further establishes stable cross-modal contrastive learning through the Gaussian-Aware Guidance module, which introduces a new perspective of leveraging pre-trained priors for stabilizing the learning of large-scale 3D representations, and provide guidance on spatial information for understanding and aligning 3DGS representations with other modalities.
@@ 4.1 Experimental Setup
Baseline. We compare UniGS with (Zeng et al., 2023; Zhang et al., 2023; 2024; Qi et al., 2023; 2024). To further understand how the 3DGS features, the pre-trained weights, and the parallel structure impact 3D representation learning, we retrain the most relevant model (Uni3D) with different settings for a fair comparison. As for the baseline model, we further report the performance of the original Uni3D model trained with point clouds and the altered version trained with the 3D location attributes (mean value of Gaussians) of 3DGS instead of point clouds for completeness. Note that 3DGS does not necessarily exist on the surface of objects, so there is a certain difference between point clouds and the 3D location attributes of 3DGS.
@@ 4.2 Comparisons to state-of-the-art
To demonstrate the effectiveness of our proposed method, we evaluate UniGS on the Text-3D retrieval, zero-shot classification, and scene understanding tasks and make comparisons with state-of-the-art methods (Zeng et al., 2023; Zhang et al., 2023). We further report the results of Uni3D (Zhang et al., 2023) trained with 3DGS for completeness. Note that CLIP2 (Zeng et al., 2023) is retrained on our collected dataset for fair comparisons.

### v2

@@ 1 Introduction
The remarkable success of 2D Image-Text pre-training through modality alignment via contrastive learning (Radford et al., 2021; Sun et al., 2023; Fang et al., 2023; Schuhmann et al., 2022; Qi et al., 2020; Changpinyo et al., 2021; Hong et al., 2021) has recently inspired a line of work pursuing 3D pre-training (Xue et al., 2023a; b; Zeng et al., 2023; Zhou et al., 2024; Liu et al., 2024; Zhang et al., 2022; Huang et al., 2023; Afham et al., 2022). Leveraging diverse large-scale 3D datasets, recent works such as  (Liu et al., 2024; Zhou et al., 2024) expand the traditional 2D task to text-image-3D pertaining by including point clouds as 3D representations, resulting in considerable improvements in 3D zero-shot/open-world object detection, classification and retrieval tasks. While point clouds serve as a natural step towards 3D representations, there are inherent limitations when using them to represent 3D objects. In particular, as illustrated in Fig. 1, point clouds consist of a discrete set of points and thus struggle to accurately capture fine-grained geometric details and surface textures of common objects, limiting the performance of 3D representation learning approaches. Moreover, there exists a noticeable gap between the discrete points and the dense 2D pixels of images, which further hinders the learning of joint multi-modal representations.
@@
However, simply retraining existing multi-modal frameworks like CLIP2 (Zeng et al., 2023) or Uni3D (Zhou et al., 2024) to leverage 3DGS as the 3D representation is not effective. As the spatial connections of 3DGS may be insufficient to capture and express objects due to the fact that 3DGS are not necessarily distributed on the surface of objects.
@@ 2 Related Work
Multi-modal pretraining via contrastive learning. Leveraging multi-modal data to pre-train modality-specific encoders via contrastive alignment has received considerable attention in recent years (Radford et al., 2021; Mu et al., 2022) due to its ability to leverage massive in-the-wild datasets of paired data. While early works have largely been focused on image-text data, promoting research on text-based image manipulation (Patashnik et al., 2021), open vocabulary object detection (Gu et al., 2021; Gao et al., 2022), language grounding (Li et al., 2022) and zero-shot segmentation (Xu et al., 2023), there has been an increasing focus on learning 3D representations lately (Xue et al., 2023a; b; Zeng et al., 2023; Zhou et al., 2024). These approaches largely follow the contrastive learning paradigm by adding a new 3D representation encoder and aligning it with the 2D and/or text modalities. However, due to the difficulty of collecting and constructing 3D representation data, current approaches leverage point clouds. More specially, PointCLIP (Zhang et al., 2022) and CLIP2Point (Huang et al., 2023) extract depth maps from point clouds to obtain image-like data that can be leveraged for contrastive pretraining. Since these depth maps lose plenty of spatial information of the original point cloud data structure, CLIP2 (Zeng et al., 2023) instead learns a 3D encoder directly on point clouds, demonstrating robustness in real-world indoor and outdoor scenarios. However, the flexibility and simplicity of point clouds come at a cost as specific shape and texture information of the object surface is lost, leading to ambiguities. In this work, we, therefore, adopt 3DGS (Kerbl et al., 2023) to replace point clouds as the 3D representation to alleviate this loss of information.
@@ 3.3 Gaussian-Aware Guidance
For better modeling and understanding of the 3D domain represented by 3DGS, we further propose a Gaussian-aware Guidance module. Specifically, this module, as highlighted in the dashed box in Fig. 3a, formulates a dual-branch 3D encoder, where the fundamental encoder $E_{fun}$ leverages the 3D ViT Encoder pre-trained on 3D point clouds from (Zhou et al., 2024) to model the low-level features $f_{fun}$ including spatial and color information of 3DGS, while the advanced encoder $E_{adv}$ additionally models the high-level features $f_{adv}$, the relationship between spatial connections and 3DGS feature. As the number of Gaussian spheres allocated to an object or scene represented by 3DGS varies, a group divider is leveraged to process the input 3DGS into a fixed number of Gaussian spheres. The CNN layers map low-dimensional raw features to a high-dimensional feature space, consistent with Uni3D. Moreover, cross-attention layers are built to extract guidance of embeddings from the fundamental encoder and improve the alignment of the 3DGS features with other domains. We denote the process as $CA$, where
@@ 3.4 Scaling Up 3DGS Backbones
Initializing UniGS. Restricted by the scale of the 3D dataset, directly pre-training each 3D backbone for specific 3D tasks leads to expensive training costs and may suffer from difficulties in convergence or overfitting. To overcome this issue, we follow Uni3D (Zhou et al., 2024) and adopt off-the-shelf pre-trained large models (Sun et al., 2023; Fang et al., 2023; Radford et al., 2021; Caron et al., 2021) with ViT-based structure in the other modalities as the initialization of the 3D backbone to transfer their rich underlying representational abilities to language-image-3D pre-training Different from Uni3D (Zhou et al., 2024), UniGS further establishes stable cross-modal contrastive learning through the Gaussian-Aware Guidance module, which introduces a new perspective of leveraging pre-trained priors for stabilizing the learning of large-scale 3D representations, and provide guidance on spatial information for understanding and aligning 3DGS representations with other modalities.
@@ 4.1 Experimental Setup
Baseline. We compare UniGS with (Zeng et al., 2023; Zhou et al., 2024; Zhang et al., 2024; Qi et al., 2023; 2024). To further understand how the 3DGS features, the pre-trained weights, and the parallel structure impact 3D representation learning, we retrain the most relevant model (Uni3D) with different settings for a fair comparison. As for the baseline model, we further report the performance of the original Uni3D model trained with point clouds and the altered version trained with the 3D location attributes (mean value of Gaussians) of 3DGS instead of point clouds for completeness. Note that 3DGS does not necessarily exist on the surface of objects, so there is a certain difference between point clouds and the 3D location attributes of 3DGS.
@@ 4.2 Comparisons to state-of-the-art
To demonstrate the effectiveness of our proposed method, we evaluate UniGS on the Text-3D retrieval, zero-shot classification, and scene understanding tasks and make comparisons with state-of-the-art methods (Zeng et al., 2023; Zhou et al., 2024). We further report the results of Uni3D (Zhou et al., 2024) trained with 3DGS for completeness. Note that CLIP2 (Zeng et al., 2023) is retrained on our collected dataset for fair comparisons.